# ACSE-Eval: Can LLMs threat model real-world cloud infrastructure?

### v1

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated promising performance in cybersecurity tasks such as vulnerability detection and code analysis [Chen et al., 2023] [Ferrag et al., 2025]. However, their ability to perform architectural threat assessments in complex cloud environments remains under explored. While effective at identifying source-level issues [Li and Shin, 2024], LLMs have yet to prove they can reason about service interactions, trust boundaries, and multi-resource configurations typical of modern cloud systems. The urgency of this research is underscored by the evolving threat landscape in cloud security. In 2024, cloud breaches have reached alarming levels, with 79% of cloud-based enterprises reporting at least one incident, and about 25% of organizations expressed uncertainty about undetected threats [Johnson and Patel, 2025, SentinelOne Research Team, 2024]. More concerning is that 82% of these breaches originated from architectural design flaws or misconfigurations, highlighting a critical gap in current security approaches.
Traditional approaches to threat modeling, exemplified by frameworks like MITRE ATT&CK [MITRE, 2025] and STRIDE [Microsoft, 2009], have been enhanced with cloud-specific considerations. However, these frameworks, while comprehensive in mapping adversarial tactics and techniques, have yet to fully integrate with LLM capabilities for automated threat assessment. The complexity of modern cloud architectures, combined with the dynamic nature of Identity and Access Management (IAM) and Zero Trust requirements [Rose et al., 2020], demands more sophisticated threat modeling capabilities that can adapt to rapidly evolving security challenges.
Recent initiatives in the LLM-enabled threat-modeling space are steps in the right direction. Elsharef et al. [2024] explores the practical applicability of an LLM assisted threat model but lacks a comprehensive evaluation procedure. Projects such as Auspex [Gupta et al., 2023] and ThreatModeling-LLM [Yang et al., 2024] aim to address this. However, they are focused on industry-specific infrastructure such as banking. Existing datasets primarily evaluate LLMs using Capture-the-Flag (CTF) challenges and exploitation tasks [Bhatt et al., 2024, Huang and Zhu, 2024, Wan et al., 2024, Kouremetis et al., 2025]. While valuable, these gamified, bounded tasks fail to reflect the complexity of real-world cloud infrastructure. Dynamic testing frameworks such as Phuong et al. [2024], Alam et al. [2024] and red teaming initiatives like DARPA AIxCC broaden this scope but still emphasize exploit generation over architectural reasoning. Moreover, many tasks risk conflating memorization with genuine security understanding. Recently, there have been efforts to benchmark general cybersecurity knowledge [Tihanyi et al., 2024, Chen et al., 2024, Liu et al., 2024], threat intelligence [Alam et al., 2024], and IaC security [Kon et al., 2024], but they do not address holistic threat modeling across cloud architectures and their corresponding IaC artifacts.
To address these limitations, we present ACSE-Eval, a dataset designed to evaluate LLMs’ threat modeling capabilities. Our contributions include: a) a curated dataset of 100 real-world AWS architecture diagrams with Diagrams-as-Code [Kwon, 2020] and Infrastructure-as-Code implementations using AWS CDK [Amazon Web Services, 2024], b) expert-generated threat models aligned with STRIDE [Shostack, 2014], ATT&CK [MITRE, 2025], and OWASP Top 10 [OWASP, 2025], c) a multi-dimensional evaluation framework assessing threat identification, vulnerability analysis, and mitigation suggestions, and d) an open-source release of the dataset and evaluation toolkit to promote further research. Spanning 100 human-annotated threat scenarios and over 2,500 hours (about 3 months) of expert effort, our dataset targets AWS, representing 31% of the cloud market [TechCrunch, 2024] and includes use cases ranging from simple web apps to multi-region, and hybrid deployments.
@@
In addition to advancing the field of cybersecurity evaluations for LLMs via the open-source dataset, our work also aims to address the following key research questions:
@@ 2.2 Dataset Overview
Figure 3: Left: Category-wise breakdown of the AWS services that are part of ACSE-Eval. Right: Distribution of OWASP Top 10 threats in the ACSE-Eval dataset.
@@
Component Relationship Context (CRC). Each threat model and IaC implementation is accompanied by a PNG architecture diagram and its corresponding Python-based diagram generation code Kwon [2020]. While our current scope focuses on textual inputs and evaluation, these architectural visualizations serve as valuable artifacts for establishing relationships between various system components. We hypothesize that providing LLMs with this component relationship context, alongside the deployment code, could enhance the quality of generated threat models. Future research could extend this work by incorporating visual (multi-modal) in-context learning techniques, potentially leading to more comprehensive and accurate threat modeling capabilities.
@@ 3.1 Implementation and Tooling
Our evaluations leverage the managed LLM APIs provided by Amazon [2023], Anthropic [2024], DeepMind [2023], and OpenAI [2023]. The comprehensive experiment incurred a cost of approximately \$500. The evaluation framework is implemented using Inspect AI, an open-source platform for LLM assessments developed by UK AI Security Institute [2024]. During our analysis, we employed specialized packages for text similarity measurements, including rouge for calculating Rouge-L scores [Lin, 2004] and sentence-transformers for computing semantic (cosine) similarity [Reimers and Gurevych, 2019]. To ensure transparency and reproducibility, the source code for both the experiment and evaluation procedures is freely available at https://github.com/ACSE-Eval/acse-eval-experiments under the MIT License.

### v2

@@ Abstract
Keywords:
LLM evaluation Automated threat-modeling Cloud security Dataset
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated promising performance in cybersecurity tasks such as vulnerability detection and code analysis [7] [10]. However, their ability to perform architectural threat assessments in complex cloud environments remains under explored. While effective at identifying source-level issues [17], LLMs have yet to prove they can reason about service interactions, trust boundaries, and multi-resource configurations typical of modern cloud systems. The urgency of this research is underscored by the evolving threat landscape in cloud security. In 2024, cloud breaches have reached alarming levels, with 79% of cloud-based enterprises reporting at least one incident, and about 25% of organizations expressed uncertainty about undetected threats [13, 26]. More concerning is that 82% of these breaches originated from architectural design flaws or misconfigurations, highlighting a critical gap in current security approaches.
Traditional approaches to threat modeling, exemplified by frameworks like MITRE ATT&CK [21] and STRIDE [20], have been enhanced with cloud-specific considerations. However, these frameworks, while comprehensive in mapping adversarial tactics and techniques, have yet to fully integrate with LLM capabilities for automated threat assessment. The complexity of modern cloud architectures, combined with the dynamic nature of Identity and Access Management (IAM) and Zero Trust requirements, demands more sophisticated threat modeling capabilities that can adapt to rapidly evolving security challenges.
@@
Recent initiatives in the LLM-enabled threat-modeling space are steps in the right direction. Elsharef et al. [9] explores the practical applicability of an LLM assisted threat model but lacks a comprehensive evaluation procedure. Projects such as Auspex [11] and ThreatModeling-LLM [32] aim to address this. However, they are focused on industry-specific infrastructure such as banking. Existing datasets primarily evaluate LLMs using Capture-the-Flag (CTF) challenges and exploitation tasks [5, 12, 31, 15]. While valuable, these gamified, bounded tasks fail to reflect the complexity of real-world cloud infrastructure. Dynamic testing frameworks [24, 1] and red teaming initiatives like DARPA AIxCC broaden this scope but still emphasize exploit generation over architectural reasoning. Moreover, many tasks risk conflating memorization with genuine security understanding. Recently, there have been efforts to benchmark general cybersecurity knowledge [29, 6, 19], threat intelligence [1], and IaC security [14], but they do not address holistic threat modeling across cloud architectures and their corresponding IaC artifacts.
To address these limitations, we present ACSE-Eval, a dataset designed to evaluate LLMs’ threat modeling capabilities. Our contributions include: a) a curated dataset of 100 real-world AWS architecture diagrams with Diagrams-as-Code [16] and Infrastructure-as-Code implementations using AWS CDK [3], b) expert-generated threat models aligned with STRIDE [27], ATT&CK [21], and OWASP Top 10 [23], c) a multi-dimensional evaluation framework assessing threat identification, vulnerability analysis, and mitigation suggestions, and d) an open-source release of the dataset and evaluation toolkit to promote further research. Spanning 100 human-annotated threat scenarios and over 2,500 hours (about 3 months) of expert effort, our dataset targets AWS, representing 31% of the cloud market [28] and includes use cases ranging from simple web apps to multi-region, and hybrid deployments.
In addition to advancing the field of cybersecurity evaluations for LLMs via the open-source dataset, our work aims to address the following key research questions:
@@ 2.1 Dataset Generation Workflow
Figure 3: Category-wise breakdown of the AWS services that are part of ACSE-Eval.
@@ 2.2 Dataset Overview
Figure 4: Distribution of OWASP Top 10 threats in the ACSE-Eval dataset.
@@
Figure 5: Heatmaps showing the relationship between threat likelihood and CIA (Confidentiality, Integrity, Availability) impact across all architectures evaluated in ACSE-Eval.
Component Relationship Context (CRC). Each threat model and IaC implementation is accompanied by a PNG architecture diagram and its corresponding Python-based diagram generation code [16]. While our current scope focuses on textual inputs and evaluation, these architectural visualizations serve as valuable artifacts for establishing relationships between various system components. We hypothesize that providing LLMs with this component relationship context, alongside the deployment code, could enhance the quality of generated threat models. Future research could extend this work by incorporating visual (multi-modal) in-context learning techniques, potentially leading to more comprehensive and accurate threat modeling capabilities.
@@ 3.1 Implementation and Tooling
Our evaluations leverage the managed LLM APIs provided by Amazon [2], Anthropic [4], Google [8], and OpenAI [22]. The comprehensive experiment incurred a cost of approximately \$500. The evaluation framework is implemented using Inspect AI, an open-source platform for LLM assessments developed by [30]. During our analysis, we employed specialized packages for text similarity measurements, including rouge for calculating Rouge-L scores [18] and sentence-transformers for computing semantic (cosine) similarity [25]. To ensure transparency and reproducibility, the source code for both the experiment and evaluation procedures is freely available at https://github.com/ACSE-Eval/acse-eval-experiments under the MIT License.